OpenStudy (anonymous):

3 (y+6) = 30 solve the equation

OpenStudy (solomonzelman):

hint: \[30= 3 \times 10\] so you can say \[\color{blue} { 3 } (y+6)=\color{blue} { 3 } \times 10\] see what you can do?

OpenStudy (anonymous):

not really im new to this type of problem

OpenStudy (solomonzelman):

You can divide both sides by 3.

OpenStudy (anonymous):

but it would be 10?

OpenStudy (solomonzelman):

yes, the right side would be equal 10, what about the left side?

OpenStudy (anonymous):

2?

OpenStudy (solomonzelman):

No, you have \[3(y+6)~~~~~which~~~is~~~same~~~as,~~~~3 \times (y+6)\]

OpenStudy (solomonzelman):

divide the left side by 3, and you get ?

OpenStudy (anonymous):

ohhh... ok well can you show me step by step with the answer that way i can know how to do this problem

OpenStudy (solomonzelman):

\[3(y+6)=30\]\[\color{red}{ \frac{\color{black} { 3(y+6) } }{3} \color{black} { = } \frac{\color{black} { 30 } }{3} } \]\[y+6=10\]\[y+6\color{red} { -6 } =10\color{red} { -6 } \]\[y=?\]

OpenStudy (anonymous):

y= 5

OpenStudy (solomonzelman):

\[10-6=?\]

OpenStudy (anonymous):

OOHHHH SOO FOUR

OpenStudy (solomonzelman):

Yeah, 4.

OpenStudy (anonymous):

THANK YOU HERES THE MEDAL

OpenStudy (solomonzelman):

You welcome!